How do you use only the number 5, 5 times in a equation to equal 17?
Answer by nyc_function(2741)

Any number raised to the zero power = 1.
Keep in mind that 5^0 = 1.
5 + 5 + 5 + 5^0 + 5^0 = 17
15 + 1 + 1 = 17
15 + 2 = 17
17 = 17
The equation is 5 + 5 + 5 + 5^0 + 5^0 = 17